Calculate the work done when a force F = 2i+3j-5k units acts on a body, producing a displacement S = 2i+4j+3k units
1
Expert's answer
2020-11-05T07:34:12-0500
"W=(2,3,-5)\cdot (2,4,3)\\W=2(2)+3(4)-5(3)=1\ J"